### v1

# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

Abstract
While 3D instance segmentation has made significant progress, current methods struggle to address realistic scenarios where new categories emerge over time with natural class imbalance. This limitation stems from existing datasets, which typically feature few well-balanced classes. Although few datasets include unbalanced class annotations, they lack the diverse incremental scenarios necessary for evaluating methods under incremental settings. Addressing these challenges requires frameworks that handle both incremental learning and class imbalance. However, existing methods for 3D incremental segmentation rely heavily on large exemplar replay, focusing only on incremental learning while neglecting class imbalance. Moreover, frequency-based tuning for balanced learning is impractical in these setups due to the lack of prior class statistics. To overcome these limitations, we propose a framework to tackle both Continual Learning and class Imbalance for 3D instance segmentation (CLIMB-3D). Our proposed approach combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module. Unlike prior methods, our framework minimizes ER usage, with KD preventing forgetting and supporting the IC module in compiling past class statistics to balance learning of rare classes during incremental updates. To evaluate our framework, we design three incremental scenarios based on class frequency, semantic similarity, and random grouping that aim to mirror real-world dynamics in 3D environments. Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline. Code is available at: https://github.com/vgthengane/CLIMB3D
1 Introduction
Figure 1: Class-incremental setup for 3D instance segmentation. As tasks progress through time, new classes are introduced incrementally. After each new task, the model should recognize both previously learned and newly introduced classes. For example, at Task-2, new classes such as Pillow, Coffee Table, and Sofa Chair are added, and the model is able to detect these three classes along with previous ones like Table, Chair, and Couch.
3D instance segmentation is an essential task in computer vision that involves identifying and segmenting individual objects in the real physical space, playing a key role in applications across graphics, robotics, and autonomous systems. Its ability to provide precise object boundaries and class labels enhances scene understanding, facilitates object manipulation, and improves perception in dynamic environments.
In recent years, a variety of methods have been proposed, including top-down approaches [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up approaches [hou20193dsis, yang2019learning], and transformer-based architectures [Schult23ICRA]. These methods have shown impressive results in traditional setups, which assume that all object classes are available during training. However, this assumption limits its applicability in real-world scenarios where new categories gradually emerge over time, often exhibiting naturally imbalanced distributions. Thus, there is a need for class-incremental learning (CIL) frameworks that not only adapt to new classes but also preserve prior knowledge, especially for rare or less frequent categories, which are more prone to catastrophic forgetting.
Most existing research in class-incremental learning focuses on 2D image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with some extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. These methods employ strategies such as exemplar replay [buzzega2020dark, rebuffi2017icarl, cha2021co2l, kamra2017deep], regularization [aljundi2018memory, li2017learning, serra2018overcoming], and knowledge distillation [douillard2020podnet, kang2022class] to preserve previously learned knowledge and mitigate catastrophic forgetting [mccloskey1989catastrophic]. Few studies have applied CIL to point clouds; however, they mostly focus on object-level classification [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. At the scene level, some works have explored 3D semantic segmentation [Yang_2023_CVPR] with incremental learning, but their performance is not as competitive as state-of-the-art methods that do not employ incremental learning, which limits their applicability. Other methods tackle open-world incremental learning [boudjoghra20243d] but rely heavily on large exemplar sets [rolnick2019experience] and often neglect the challenge of class imbalance.
To address this, we propose CLIMB-3D, a unified framework that combines Exemplar Replay (ER), Knowledge Distillation (KD), and a novel Imbalance Correction (IC) module to tackle Continual Learning for Imbalance 3D instance segmentation in indoor environments, as shown in Figure 1. Our framework operates as follows: ER stores a subset of representative samples from previous stages, allowing for effective replay during new task learning. KD transfers knowledge by retaining a copy of the model from the previous task, thereby mitigating forgetting. The IC module is specifically designed to reduce forgetting of rare classes by leveraging the frequency of object occurrence. However, since we do not have access to the previous task data or statistics during incremental phases, we instead compile these statistics from the previous model used by KD to generate weight for earlier categories. These weights are used to favor both frequent and rare classes, ensuring a balanced learning and mitigating forgetting.
To evaluate CLIMB-3D in a realistic, incremental learning setup, we create three benchmark scenarios using the ScanNet200 dataset [rozenberszki2022language], which features 200 classes with natural class imbalances. These scenarios are designed to reflect real-world conditions where new categories emerge gradually and follow inherent class imbalances. These are based on 1 frequency of object occurrence, 2 semantic similarity between the object, and 3 random grouping. Our experiments demonstrate that CLIMB-3D significantly improves performance by effectively mitigating the forgetting of previous tasks compared to earlier exploration in class incremental 3D segmentation.
In summary, our contributions are as follows:
  1. 1.
    We propose a new problem setting for imbalanced class incremental 3D segmentation, along with a simple yet effective method to address this challenge by minimizing catastrophic forgetting and balancing the learning process.
  2. 2.
    To benchmark this setting, we design three scenarios aimed at simulating real-world conditions where objects emerge continuously with natural class imbalance.
  3. 3.
    Experimental results show that our proposed framework achieves state-of-the-art performance, with an increase of up to 16.76% in mAP compared to the baseline.
2 Related Work
This section reviews the current literature on 3D instance segmentation and incremental learning methods, including the limited work addressing incremental learning for 3D scene-level tasks.
2.1 3D Instance Segmentation
Various approaches have been proposed for 3D instance segmentation. One common approach adopts a bottom-up pipeline, in which an embedding in the latent space is learned to facilitate the clustering of object points [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance]. These methods are also known as grouping-based or clustering-based methods. Other methods use a top-down approach, also known as proposal-based methods, where 3D bounding boxes are first detected, then the object region is segmented within the box [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d]. Recently, the transformer architecture [vaswani2017attention] has also been used for the task of 3D instance segmentation [Schult23ICRA, sun2022superpoint], motivated by work in 2D [cheng2022masked, cheng2021per]. While these methods propose various models for improving the quality of the object segments, they rely on the availability of annotations for all object categories. On the other hand, we target learning in a progressive manner, in which new semantic annotation is provided and past data is inaccessible.
In order to reduce the annotation cost for 3D instance segmentation, various methods propose weakly supervised alternatives to methods that use dense annotations [xie2020pointcontrast, hou2021exploring, chibane2022box2mask]. While these methods improve the ability to learn from a small set of annotated examples, they rely on a fixed set of semantic labels, so they are prone to catastrophic forgetting in an incremental setting.
2.2 Incremental Learning
Incremental, lifelong, or continual learning methods aim to train a machine learning model sequentially to avoid “catastrophic forgetting” which is caused by training the model on a set of data and later training on another set of data. There are several methods have been proposed for this paradigm, these methods can be divided into three categories: (i) Model Regularization [kirkpatrick2017overcoming, aljundi2018memory, li2017learning] methods limit the plasticity of model parameters to avoid catastrophic forgetting of previous tasks. These methods include weight regularization such as EWC [serra2018overcoming] and function regularization such as knowledge distillation [hinton2015distilling]. (ii) Exemplar replay approaches either create a subset of the past task data or generate samples using generative models to avoid privacy concerns and save those in memory to replay while learning new tasks. This method is effective in more challenging settings and datasets [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l]. (iii) Dynamic network expansion-based method learns a new task by either dynamically expanding the model [rusu2016progressive, li2019learn, zhao2022deep] or by creating a subset of the model [zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random] to learn to cater for a new task.
Recent approaches to 3D class-incremental segmentation, such as [Yang_2023_CVPR] and [su2024balanced], have made some initial contributions. However, these methods often fall short in performance as they do not leverage state-of-the-art 3D segmentation models and are primarily focused on semantic segmentation, while our work emphasizes object-level instance segmentation. Kontogianni et al. [kontogianni2024continual] propose a general online continual learning framework and evaluate it on 3D dataset segmentation. Similarly, [boudjoghra20243d] addresses the open-world 3D incremental learning problem but relies heavily on an extensive memory buffer. In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
3 Preliminaries
3.1 Transformer-based Segmentation
We adopt a transformer-based instance segmentation method based on Mask3D [Schult23ICRA]. Specifically, transformer-based segmentation model $\Phi$ is employed for point cloud instance segmentation. Given an input point cloud $p$, the model predicts $\hat{y}=\{(\hat{m_{j}},\hat{c_{j}})\}_{j=1}^{J}$, which consists of mask predictions and class probabilities for each instance. The segmentation process begins by quantizing the input point cloud $p$ into voxels $V$, creating voxelized representations of size $\mathbb{R}^{M_{0}\times 3}$. Each voxel is assigned an average RGB color computed from the points within that voxel, serving as its initial feature representation. The feature backbone network generates a high-resolution output feature map $\mathbf{F}_{0}\in\mathbb{R}^{M_{0}}$. Additionally, intermediate feature maps are extracted from the decoder layers of the backbone network. For each intermediate feature map ($r\geq 0$), a set of $K_{r}$ voxels is selected, and their features are linearly projected to a fixed dimension $D$, yielding feature maps $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$.
The Transformer decoder initiates with a set of K instance queries and iteratively improves them using L Transformer decoder layers. These layers employ cross-attention to refine the instance queries, incorporating information from point cloud features. The decoder attends to a specific feature map obtained from the corresponding feature backbone layer at each layer, employing conventional cross-attention mechanisms. This process enables the decoder to reason at the instance level through self-attention, resulting in the generation of accurate and contextually relevant instance queries tailored to the specific scene.
To achieve this, the voxel features $\mathbf{F}_{r}\in\mathbb{R}^{M_{r}\times D}$ are transformed into sets of keys $\mathbf{K}\in\mathbb{R}^{M_{r}\times D}$ and values $\mathbf{V}\in\mathbb{R}^{M_{r}\times D}$ through linear projection. The instance queries $\mathbf{Z}$ are also projected to create the queries $\mathbf{Q}$. This enables cross-attention, allowing the queries to gather relevant information from the voxel features. Following cross-attention, a self-attention step occurs among the queries, facilitating information exchange and refinement. The learned queries are then used to make $K$ class and mask predictions, which are matched with ground truth labels through bipartite matching, resulting in $\hat{y}=\{(\hat{m}_{j},\hat{c}_{j})\}_{j=1}^{J}$. The model is optimized based on the ground truth label, mask, and class predictions:
$\mathcal{L}_{\mathrm{Seg}}(y_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m_{j},\hat{m}_{j})+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(c_{j},\hat{c}_{j})$ (1)
where, mask loss $\mathcal{L}_{\mathrm{mask}}=\lambda_{\mathrm{ce}}\mathcal{L}_{\mathrm{ce}}(y_{j},\hat{y}_{j})+\lambda_{\mathrm{dice}}\mathcal{L}_{\mathrm{dice}}(y_{j},\hat{y}_{j})$ and $\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}$ is classification loss.
The traditional setup assumes all categories are available and well-balanced during training. However, in scenarios where only a subset of categories is present, training the model in multiple phases is required. Unfortunately, such multi-phase models often suffer from forgetting previous tasks. To address this issue, we employ incremental learning strategies which will be discussed in the next section.
4 Methodology
In this section, we begin by formulating the incremental 3D instance segmentation problem and introduce our proposed method to address it using a transformer-based model. We then detail the design of incremental scenarios, outlining the motivation and considerations behind each one.
Figure 2: CLIMB-3D: At phase $t$ (for $t>1$), we create a copy of the model from the previous phase, $\Phi^{\mathrm{T}}$ (the teacher model), and designate it as $\Phi^{\mathrm{S}}$ (the student model). Both $\Phi^{\mathrm{T}}$ and $\Phi^{\mathrm{S}}$ process the input point cloud simultaneously, producing predictions $\hat{y}_{\mathrm{pseudo}}$ and $\hat{y}$, respectively. To prevent the model from forgetting less frequent categories from previous tasks, we balance the predictions from $\Phi^{\mathrm{T}}$, resulting in a balanced output $\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}$. The concatenated vector of ground truth labels $[y,\hat{y}_{\mathrm{pseudo}},\hat{y}^{\mathrm{(w)}}_{\mathrm{pseudo}}]$ is then compared with the predicted labels $\hat{y}$. A loss function is applied based on this comparison, enabling the student model to learn from the differences between the predicted and the concatenated pseudo-ground truth labels
4.1 Problem Formulation
The objective of 3D point cloud instance segmentation is to accurately identify and segment individual instances of objects within a given point cloud. Mathematically, the training dataset is represented as $\mathcal{D}=(\mathcal{P},\mathcal{Y})=\{(p_{i},y_{i})\}_{i=1}^{N}$, where $N$ is the total number of samples. Each sample consists of a colored point cloud $p_{i}\in\mathbb{R}^{M\times 6}$ of size $M$, where the point coordinates and color values are represented as $\{x,y,z,r,g,b\}$. The corresponding annotations are denoted as $y_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J}$, where $m_{i,j}$ represents the instance mask for the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ denotes the semantic label of the object category to which the instance belongs for the $i$-th point cloud. Here, $J$ represents the total number of instances in the $i$-th point cloud, and $C$ indicates the number of distinct object categories. During the learning process, the model $\Phi$ will process this dataset and output predictions $\hat{y}_{i,j}=(\hat{m}_{i,j},\hat{c}_{i,j})$, where $\hat{m}_{i,j}$ represents the predicted instance mask and $\hat{c}_{i,j}$ denotes the predicted semantic label for the $j$-th instance in the $i$-th point cloud.
To adapt the dataset to an incremental learning setting, we partition the object categories $\mathcal{C}$ into $T$ subsets, denoted as $\mathcal{C}=\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{T}$. Each phase $t\in\{1,\dots,T\}$ is associated with a specific subset $\mathcal{C}^{t}$, and its corresponding dataset $\mathcal{D}^{t}$ which only contains annotations for objects belonging to the corresponding subset. Formally, during the $t$-th phase of training, the dataset $\mathcal{D}^{t}=(\mathcal{P},\mathcal{Y}^{t})=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ is defined, where $\mathcal{P}$ represents the point clouds shared across all phases, and $\mathcal{Y}^{t}$ contains annotations exclusively for objects belonging to the class subset $\mathcal{C}^{t}$. It is important to note that the 3D scenes within each phase can contain objects of any type from the entire object category set $\mathcal{C}$, but only the object belonging to $\mathcal{C}^{t}$ are annotated during that specific phase. After training for phase $t$ completes, the model is evaluated on a validation set containing the union of classes up to task $t$ (i.e., $\mathcal{C}^{1}\cup\cdots\cup\mathcal{C}^{t}$). Training progresses to the next phase, $t+1$, where the model $\Phi$ observes the same set of 3D scenes $\mathcal{P}$ but with annotations for different object types belonging to the subset $\mathcal{C}^{t+1}$. This incremental training approach allows the model to gradually learn and adapt to new object categories over multiple phases.
4.2 CLIMB-3D
In our proposed framework (Figure 2), the incremental instance segmentation model undergoes phased training, as described in Sec. 4.1, where carefully designed subsets of the dataset are introduced to handle various real-world scenarios discussed in Sec. 4.3. Formally, at phase $t$, when the model $\Phi^{t}$ is introduced with input data $\mathcal{D}^{t}=\{(p_{i},y_{i}^{t})\}_{i=1}^{N}$ and trained using Eq. 1, a common issue arises where it tends to forget the knowledge acquired in the previous phase, leading to catastrophic forgetting [mccloskey1989catastrophic]. To address this, we first draw inspiration from techniques developed in the 2D domain [rebuffi2017icarl, li2017learning] and recent 3D semantic segmentation [su2024balanced, Yang_2023_CVPR], and adapt them for our setting. However, we observe that these adaptations alone fall short of achieving the desired performance levels; therefore, we propose a teacher-student knowledge distillation framework to effectively retain previously learned knowledge. Additionally, we incorporate an imbalance correction module to handle the challenge of less frequent classes from earlier tasks.
Exemplar Replay (ER). Inspired by the approach proposed by Buzzega et al. [buzzega2020dark], ER methods alleviate the issue of limited exposure to previous task data during training. By selectively storing a small subset of exemplars $\mathcal{E}_{t}$ from previous phases, the model can learn from both the current task data $\mathcal{D}_{t}$ and the replayed exemplars $\mathcal{E}_{1:t-1}$. This results in a combined dataset $\mathcal{D}_{t}\cup\mathcal{E}_{1:t-1}$, where $\mathcal{E}_{1:t-1}$ represents the exemplar memory formed by the union of all previous exemplar sets $\mathcal{E}_{1:t-1}=\mathcal{E}_{1}\cup~{}\dots~{}\cup\mathcal{E}_{t-1}$. The model undergoes a full iteration on $\mathcal{D}_{t}$ before replaying the exemplars.
Some previous 3D approaches adopt this strategy to retain knowledge but rely on a large exemplar set [boudjoghra20243d], which is often impractical in real-world scenarios. To address this, we choose a smaller exemplar set, creating a more challenging setup that requires the model to effectively manage and retain knowledge with limited resources, thereby testing its robustness and adaptability in practical applications.
Knowledge Distillation (KD) Module. In our incremental learning approach, we utilize a Knowledge Distillation (KD) module that incorporates a teacher-student framework, maintaining a copy of the previously trained model while learning the current task. For $t>1$, at the beginning of each training stage, the current model $\Phi^{t}$ is initialized as $\Phi^{t}\leftarrow\Phi^{t-1}$, where $\Phi^{t-1}$ represents the model trained in the previous phase. As the $\Phi^{t}$ is trained on the previous stage dataset, it holds information about the previous set of classes. Hence, we make use of this model to help retain previous knowledge while learning the current task.
When presented with a new training point-cloud and label pair $(p,y^{t})$, the output of the previous model is calculated as $\hat{y}_{\mathrm{pseudo}}=\Phi^{\mathrm{t-1}}(p)$, and a combined loss function is minimized. This combined loss comprises the $\mathcal{L}_{\mathrm{Seg}}(\hat{y}^{t},y^{t})$ loss, which measures the discrepancy between the predicted and ground truth labels, and the knowledge distillation loss $\mathcal{L}_{\mathrm{KD}}$, which encourages the similarity between the predictions of the current model and the previous model.
$\mathcal{L}_{\mathrm{KD}}(\hat{y}^{t},\hat{y}_{\mathrm{pseudo}})$ $=\mathcal{L}_{\mathrm{mask}}(\hat{m}^{t}_{j},\hat{m}_{j,\mathrm{pseudo}})$
$\quad\quad\quad\quad+\lambda_{\mathrm{cls}}\mathcal{L}_{\mathrm{cls}}(\hat{c}^{t}_{j},\hat{c}_{j,\mathrm{pseudo}})$ (2)
However, as pointed out in previous works on object detection in 2D [liu2023continual], Section 4.2 is often biased towards the background classes, as the model tends to predict the background for most instances. Similarly, we propose selecting the top $K$ most confident predictions from the previous teacher model $\Phi^{t-1}$ and combining them with the ground truth labels, which then serve as pseudo-labels. By extracting the top $K$ confident samples from the output of $\Phi^{t-1}$ and combining them with the ground truth labels, the augmented label set becomes $y^{\prime}=[y^{t},\hat{y}_{\mathrm{pseudo}}^{K}]$. This augmented label set is then used to optimize the current model $\Phi$ with the $\mathcal{L}_{\mathrm{Seg}}$ loss function from Equation 1.
Imbalance Correction (IC) Module. Although retaining a few samples from previous tasks and selecting the most confident predictions from the previous model, $\Phi^{t-1}$, helps preserve information from prior tasks while learning new ones, we observed that this approach does not adequately address the class imbalance. Our analysis reveals that the most confident predictions from $\Phi^{t-1}$ are largely associated with the most frequent object categories, causing the model to forget less common classes. This issue can be mitigated by re-weighting the predictions based on the frequency of observed categories [islam2021class].
At task $t$, we only have access to the data and statistics of the current task; the previous task’s dataset and statistics are unavailable. To incorporate balancing elements despite the absence of previous stage statistics, we propose leveraging pseudo prediction statistics from $\Phi^{t-1}$. During each iteration, we use $\Phi^{t-1}$ to generate pseudo labels and accumulate class frequency statistics for prior tasks throughout the current epoch training. At the end of each epoch, we combine the statistics of observed scene classes and predicted pseudo-classes, calculating the frequency $\mathbf{f}$ of all classes seen so far.
Formally, for each category $c$, we assign a weight $\mathbf{w}_{c}$ inversely proportional to its observed frequency in the predictions of $\Phi^{t-1}$ and the current dataset. The weight $\mathbf{w}_{c}$ is defined as: $\mathbf{w}_{c}=\frac{1}{\mathbf{f}(c)+\epsilon}$, where $\epsilon$ is a small constant to avoid division by zero. In the next epoch, predictions from $\Phi^{t-1}$ are re-weighted using $\mathbf{w}_{c}$, creating an adjusted high-confidence pseudo label set for less frequent categories: $\hat{y}_{\mathrm{pseudo}}^{(w)}=\mathbf{w}_{c}\cdot\hat{y}_{\mathrm{pseudo}}$. This re-weighting occurs at each epoch, allowing $\Phi^{t-1}$ to yield a broader set of less frequent classes. To ensure the model encounters both high-confidence and less frequent classes, we select the top $K$ high-confidence predictions both before and after re-weighting. The resulting augmented label space, which combines ground truth labels, original pseudo labels, and re-weighted pseudo labels, is given by: $y^{\prime}=[y,\hat{y}_{\mathrm{pseudo}}^{K},y_{\mathrm{pseudo}}^{(w),K}]$.
As this weighting scheme is applied only to previous model predictions, we further tune the current model to favor less frequent classes by incorporating the same weights $\mathbf{w}_{c}$ into the classification loss of Equation 1. The adjusted segmentation loss becomes:
$\mathcal{L}_{\mathrm{3DIS}}(y^{\prime}_{j},\hat{y}_{j})=\mathcal{L}_{\mathrm{mask}}(m^{\prime}_{j},\hat{m}_{j})+\mathbf{w}^{\prime}_{c}\mathcal{L}_{\mathrm{cls}}(c^{\prime}_{j},\hat{c}_{j}),$ (3)
where $\mathbf{w}^{\prime}_{c}=\mathbf{w}_{c}\cdot\lambda_{\mathrm{cls}}$ represents the adjusted category weights. By re-weighting and augmenting both the label space and the loss function, our IC module ensures that both the current and previous models encounter pseudo labels spanning the long-tail distribution, addressing class imbalance and enabling more balanced learning.
4.3 Designing Incremental Scenarios
Figure 3: Incremental scenarios are grouped based on frequency of occurrence, semantic similarity, and random clustering. Different color clouds ( ,  ,  ) represent tasks in each scenario, while various shapes represent object categories, and   denotes the background. Left: Tasks are organized based on the frequency of object categories. Middle: Tasks are grouped by semantic similarity, where objects with similar shapes (e.g., circles, plus signs, and triangles) denote semantically similar classes. Right: In this fully random scenario, tasks may contain a mix of semantically similar, more frequent, or less frequent classes.
While conventional incremental learning methods have numerous practical applications, they often assume an equal distribution of samples, which does not reflect real-world conditions. In practice, the number of object categories, $\mathcal{C}$, is typically large, with significant variability in category occurrence, shape, structure, and size. With these attributes in mind, we design three incremental learning scenarios, each addressing distinct aspects of real-world conditions, the design is highlighted in Figure 3.
1
Frequency Scenarios (Split_A). This scenario acknowledges that datasets are often labeled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rarer categories.
2
Semantic Scenarios (Split_B). In real-world environments, objects may exhibit similarities in appearance, and then moved to different environments, the model may encounter new objects that does not share similar semantic characteristics with previously seen categories. To address this, we introduce the Split_B  scenario. It involves grouping categories based on their semantic labels and incrementally training the model to handle these groups. This allows the model to generalize its knowledge across semantically similar categories, facilitating adaptation to new objects with similar characteristics. Unlike the Split_A  scenario, this scenario may include both frequent and infrequent categories within the same task.
3
Random Scenarios (Split_C). In some cases, data labeling is based on the availability of objects rather than specific criteria. To account for this scenario, we design the Split_C  scenario. This scenario represents a completely random setting where each task can have any class, leading to varying degrees of class imbalance. By exposing the model to such diverse and imbalanced scenarios, we aim to enhance its ability to handle real-world situations where the availability of labeled data is unpredictable.
By designing these three incremental learning scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
5 Experiments
We evaluate our method across three scenarios specifically designed to capture the complexities of real-world settings, where new categories emerge incrementally alongside class imbalance, as discussed in Section 4.3. Experimental results demonstrate that our approach effectively handles the gradual introduction of new classes and mitigates the impact of class imbalance within these scenarios. The following sections detail the datasets, evaluation metrics, incremental scenarios, and implementation procedures, followed by a thorough analysis and comparison of the results.
5.1 Experimental Setup
Datasets. We evaluate our method using the ScanNet200 dataset [rozenberszki2022language], which includes 200 object categories and exhibits inherent class imbalance, making it well-suited for simulating and evaluating real-world scenarios. Additionally, we benchmark our method against existing incremental learning approaches using the original ScanNet dataset [dai2017scannet] in a semantic segmentation setting. We follow the standard training and validation splits as defined in prior works.
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3D instance segmentation that provides a comprehensive measure of segmentation quality, accounting for both precision and recall. For comparison with existing semantic incremental learning approaches, we report mean Intersection over Union (mIoU), which measures the overlap between predicted and ground truth instances, offering a detailed evaluation of segmentation accuracy. To assess the model’s ability to mitigate catastrophic forgetting in continual learning scenarios, we use the Forgetting Percentage Points (FPP), as defined in [liu2023continual]. This metric quantifies performance degradation by measuring the accuracy drop between the initial and final training phases, on the categories observed in the first training phase.
Incremental Scenarios. As discussed in Section 4.3, we design three incremental scenarios: Split_A, Split_B, and Split_C, each consisting of three tasks, which are grouped based on object occurrence frequency, semantic similarity, and random grouping, respectively. In Split_A, the frequency of object categories progressively decreases through the tasks. This scenario follows the head, common, and tail splits present in the ScanNet200 dataset, with class distributions of 66-68-66 in each split. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, the classes are shuffled and split into three sets, resulting in 67-67-66 categories per split. These scenarios allow for a comprehensive evaluation of our approach under varying conditions, facilitating a deeper understanding of its performance and generalization in diverse real-world settings.
Implementation Details. We utilize the transformer-based model for 3D instance segmentation proposed in [Schult23ICRA], designed to iteratively attend to hierarchical feature representations. The model processes the 4 coarsest levels of a ResNet-based U-Net backbone across three iterations, progressively refining from coarse to fine, resulting in $L=12$ transformer decoder layers. Each transformer decoder layer shares weights across iterations and consists of a standard transformer layer utilizing self-attention and masked cross-attention mechanisms. The feature backbone employed is Minkowski Res16UNet34C [choy20194d].
Training Details. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA]. For joint training (Row 1, Table 4), the model is trained for 600 epochs using the AdamW optimizer [loshchilov2017decoupled] with a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, we retain the same hyperparameters, adjusting only the number of epochs, and we use a memory buffer size of 50 scenes. Training is divided into three phases, introducing one split per phase across the three designed scenarios. After each phase, the resulting model is evaluated on all classes encountered up to that point.
5.2 Results and Discussion
Table 1: Comparison between the baseline and proposed method with mAP25, mAP50, and mAP, which is after training for all the stages. We also report the FPP metric.
Scenarios Methods Average Precision $\uparrow$ FPP $\downarrow$
mAP25 mAP50 mAP mAP25 mAP50
Baseline 16.46 14.29 10.44 51.30 46.82
Split_A CLIMB-3D 35.69 31.05 22.72 3.44 2.63
Baseline 17.22 15.07 10.93 46.27 42.1
Split_B CLIMB-3D 35.48 31.56 23.69 8.00 5.51
Baseline 25.65 21.08 14.85 31.68 28.84
Split_C CLIMB-3D 31.59 26.78 18.93 9.10 7.89
To evaluate our proposed method, we conduct a comparative analysis using exemplar replay (ER) for instance segmentation and [Yang_2023_CVPR] for semantic segmentation as baselines. As shown in Table 1, our method, which integrates exemplar replay, knowledge distillation, and an imbalance correction (IC) module, achieves notable improvements over the baseline in terms of mAP and FPP. Specifically, in the Split_A  scenario, our approach significantly enhances overall performance. We observe an improvement of 19.23%, 31.05%, and 12.28% for $\text{mAP}_{50}$, $\text{mAP}_{25}$, and overall mAP, respectively, while reducing forgetting by 47.86% as measured by $\text{mAP}_{50}$. For the Split_B  scenario as well, CLIMB-3D demonstrates a consistent performance boost over the baseline, significantly improving mAP and reducing forgetting, which is lowered to 5.52% compared to 46.21% in the baseline for AP50. Likewise, in the Split_C  scenario, CLIMB-3D enhances both learning efficiency and forgetting reduction, achieving a performance of 26.78% in $\text{mAP}_{50}$  and reducing forgetting by 20.95% compared to baseline. These results across scenarios underscore the effectiveness of our approach.
Table 2: Comparison with previous method on semantic segmentation on ScanNet V2 dataset. We report the mIoU metric for the evaluated models.
Methods Phase=1 Phase=2 All
EWC [serra2018overcoming] 17.75 13.22 16.62
LwF [li2017learning] 30.38 13.37 26.13
Yang et al. [Yang_2023_CVPR] 34.16 13.43 28.98
CLIMB-3D (Ours) 69.39 32.56 59.38
Although our method focuses on segmenting individual objects (instance segmentation), we also demonstrate its performance in semantic segmentation by presenting a comparative analysis with existing methods for class-incremental semantic segmentation on the ScanNet V2 dataset (Table 2). Using our predicted labels, we assign each point the label corresponding to the highest confidence mask and exclude background labels (floor and wall), as these are not part of the object-level segmentation. Following the dataset splits established by [Yang_2023_CVPR], we report results for both training phases. Our proposed method achieves a substantial improvement over prior methods, with a gain of 35.23% in Phase 1 and approximately 19.1% in Phase 2. Overall, our method reaches a mIoU of 59.38%, significantly outperforming previous baselines, which achieve a lower mIoU of around 30%.
Table 3: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios. Each scenario is trained in three phases ($\mathtt{phase}=1,2,3$) by introducing a single split s at a time. The results highlighted in orange are with the proposed method, and the best results for each scenario are in bold.
Scenarios Methods phase=1 phase=2 phase=3
s1 s1 s2 All s1 s2 s3 All
Baseline 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28
Split_A CLIMB-3D 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05
Baseline 51.57 13.32 42.21 24.53 9.55 12.45 26.78 15.07
Split_B CLIMB-3D 51.57 46.74 37.45 43.13 46.06 15.95 26.68 31.56
Baseline 36.40 7.74 37.62 22.32 7.55 15.96 40.41 21.08
Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
We extend the analysis from Table 1 to Table 3 to highlight the impact of our proposed method on individual splits across various scenarios. The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline. For Split_A , our model shows improvement throughout the phase. In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline. The Split_B  scenario, while more complex than Split_A, achieves comparable results due to semantic similarity among classes within the same task. In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information. After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline. In the Split_C  scenario, the first-stage model struggles due to the increased complexity introduced by random grouping. In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge. Conversely, our method balances new task learning with the retention of earlier information. By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits. Overall, our proposed method improves mAP by 5.6%.
5.3 Ablation
To assess the effectiveness of each component in our proposed framework, we perform an ablation study. Initially, we establish an upper-bound performance by jointly training the model on the complete dataset using a transformer-based architecture, such as Mask3D [Schult23ICRA], referred to as the Oracle. For the incremental learning setup, we generate training splits according to the scenarios outlined earlier. In this study, we first train the model naively across phases and then sequentially integrate each module to evaluate its individual contribution to performance. Table 4 summarizes the results for the Split_A  scenario, using both the $\text{mAP}_{50}$  and FPP metrics.
Table 4: Ablation study results illustrating the impact of exemplar replay, knowledge distillation, and imbalance correction modules in a three-phase training setup. Each split, representing the subset of data introduced at each phase (p), is labeled as ‘s’ followed by the phase number. The final column, ‘All’, in each phase reports performance across all classes encountered up to that phase. Joint training results (Oracle) are highlighted in gray, while results with all modules combined are marked in orange. The best-performing results are shown in bold.
Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
s1 s1 s2 All s1 s2 s3 All
1. Oracle - - - - 55.14 30.77 25.30 37.68 -
2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
4. + KD 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
5. + IC 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
Naïve Training. In the naive incremental training setup, where no dedicated modules are incorporated, the model learns the current task but suffers from catastrophic forgetting of the previously learned tasks, as expected. This behavior is evident in row 2, where, upon transitioning to phase 2, the model entirely forgets the classes learned during phase 1. A similar trend is observed in phase 3, and this pattern is also reflected in the FPP metric.
Effect of Exemplar Replay. To mitigate catastrophic forgetting, we incorporate exemplar replay, which stores and replays examples from previous tasks. As shown in row 3, exemplar replay improves average precision by 18.5% for s1 in phase 2 and 10.38% in phase 3. It also reduces forgetting for s2 by 9.43% in phase 3, while slightly improving learning on the current task. However, substantial forgetting persists, as reflected in the FPP metric, highlighting the limitations of exemplar replay alone.
Effect of Knowledge Distillation. The addition of knowledge distillation (KD), which retains a copy of the model from previous tasks, facilitates the preservation of past task knowledge while enabling forward knowledge transfer. As shown in row 4, KD considerably reduces forgetting and boosts performance on the current task. Specifically, for s1, KD improves $\text{mAP}_{50}$ by 31.49% in phase 2 and by 39.40% in phase 3, compared to exemplar replay. Overall, KD leads to a 15.00% increase in performance while reducing forgetting by 39.40% after all tasks have been learned.
Effect of Imbalance Correction. The imbalance correction module addresses the class imbalance in the dataset by re-weighting the teacher model’s predictions during KD based on class frequency. As highlighted in row 5 of Table 4, this addition further improves performance. Specifically, for s1, imbalance correction reduces forgetting by 4.67% and 4.41% in phases 2 and 3, respectively, compared to the results without this module (row 4). For s2, while a slight decrease in current task performance is observed in phase 2, this is likely due to the module’s prioritization of mitigating forgetting less frequent classes in previous tasks. In phase 3, performance on s2 and s3 improves. Overall, imbalance correction significantly reduces forgetting, achieving improvements of 4.41% and 43.81% over KD and exemplar replay, respectively.
6 Conclusion
We address the challenge of class-incremental 3D instance segmentation with class imbalance. We propose an innovative approach that integrates a memory-efficient exemplar replay buffer, knowledge distillation, and a novel imbalance correction module. This framework mitigates the forgetting of rare classes during incremental learning by accounting for the frequency of object occurrences. To enable comprehensive evaluation, we design three incremental learning scenarios, each comprising three phases that reflect real-world dynamics. Our experimental results demonstrate that the proposed framework significantly enhances the learning of new classes while reducing forgetting of previously learned ones. The carefully designed scenarios and framework not only offer a strong baseline but also provide a clear benchmark for future research, laying a foundation for more advanced techniques in class-incremental learning.
Supplementary Material
In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix A. Next, we provide detailed split information for all scenarios, based on class names, in Appendix B. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix C.
Appendix Appendix A Evaluation on Rare Categories
The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
paper towel dispenser 2 73.10 74.90
recycling bin 3 55.80 60.50
ladder 5 53.90 57.10
trash bin 7 31.50 57.30
bulletin board 8 23.30 38.20
shelf 11 48.00 50.50
dresser 12 44.00 55.80
copier 12 93.30 94.50
object 12 3.10 3.30
stairs 13 51.70 67.70
bathtub 16 80.30 86.60
oven 16 1.50 3.30
divider 18 36.40 45.00
column 20 57.30 75.00
Average - 46.66 54.98
The results, shown in Table 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied. Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
piano 1 7.10 59.40
bucket 1 21.10 31.50
laundry basket 1 3.80 17.40
dresser 2 55.00 55.40
paper towel dispenser 2 32.50 35.50
cup 2 24.70 30.30
bar 2 35.40 39.50
divider 2 28.60 42.40
case of water bottles 2 0.00 1.70
shower 3 0.00 45.50
mirror 8 56.00 68.80
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
bathroom counter 3 3.90 20.30
ottoman 4 32.60 36.20
storage bin 3 5.10 10.50
dishwasher 3 47.40 66.20
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
sofa chair 6 14.10 43.50
file cabinet 6 49.20 57.60
tv stand 7 67.70 68.60
mirror 8 56.00 68.80
blackboard 8 57.10 82.80
clothes dryer 9 1.70 3.20
toaster 9 0.10 25.90
wardrobe 10 22.80 58.80
jacket 10 1.20 4.10
Average - 32.08 44.21
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels. The three scenarios Split_A, Split_A, and Split_C  are each divided into three tasks: Task 1, Task 2, and Task 3.
Split_A Split_B Split_C
Task 1 Task 2 Task 3 Task 1 Task 2 Task 3 Task 1 Task 2 Task 3
chair wall pillow tv stand cushion paper broom fan rack
table floor picture curtain end table plate towel stove music stand
couch door book blinds dining table soap dispenser fireplace tv bed
desk cabinet box shower curtain keyboard bucket blanket dustpan soap dish
office chair shelf lamp bookshelf bag clock dining table sink closet door
bed window towel tv toilet paper guitar shelf toaster basket
sink bookshelf clothes kitchen cabinet printer toilet paper holder rail doorframe chair
toilet curtain cushion pillow blanket speaker bathroom counter wall toilet paper
monitor kitchen cabinet plant lamp microwave cup plunger mattress ball
armchair counter bag dresser shoe paper towel roll bin stand monitor
coffee table ceiling backpack monitor computer tower bar armchair copier bathroom cabinet
refrigerator whiteboard toilet paper object bottle toaster trash bin ironing board shoe
tv shower curtain blanket ceiling bin ironing board dishwasher radiator blackboard
nightstand closet shoe board ottoman soap dish lamp keyboard vent
dresser computer tower bottle stove bench toilet paper dispenser projector toaster oven bag
stool board basket closet wall basket fire extinguisher potted plant paper bag paper
bathtub mirror fan couch fan ball coat rack structure projector screen
end table shower paper office chair laptop hat end table picture pillar
dining table blinds person kitchen counter person shower curtain rod tissue box purse range hood
keyboard rack plate shower paper towel dispenser paper cutter stairs tray coffee maker
printer blackboard container closet oven tray fire extinguisher couch handicap bar
tv stand rail soap dispenser doorframe rack toaster oven case of water bottles telephone pillow
trash can radiator telephone sofa chair piano mouse water bottle shower curtain rod decoration
stairs wardrobe bucket mailbox suitcase toilet seat cover dispenser ledge trash can printer
microwave column clock nightstand rail storage container shower head closet wall object
stove ladder stand washing machine container scale guitar case cart mirror
bin bathroom stall light picture telephone tissue box kitchen cabinet hat ottoman
ottoman shower wall pipe book stand light switch poster paper cutter water pitcher
bench mat guitar sink light crate candle storage organizer refrigerator
washing machine windowsill toilet paper holder recycling bin laundry basket power outlet bowl vacuum cleaner divider
copier bulletin board speaker table pipe sign plate mouse toilet
sofa chair doorframe bicycle backpack seat projector person paper towel roll washing machine
file cabinet shower curtain rod cup shower wall column candle storage bin laundry detergent mat
laptop paper cutter jacket toilet bicycle plunger microwave calendar scale
paper towel dispenser shower door paper towel roll copier ladder stuffed animal office chair wardrobe dresser
oven pillar machine counter jacket headphones clothes dryer whiteboard bookshelf
piano ledge soap dish stool storage bin broom headphones laundry basket tv stand
suitcase light switch fire extinguisher refrigerator coffee maker guitar case toilet seat cover dispenser shower door closet rod
recycling bin closet door ball window dishwasher dustpan bathroom stall door curtain plant
laundry basket shower floor hat file cabinet machine hair dryer speaker folded chair counter
clothes dryer projector screen water cooler chair mat water bottle keyboard piano suitcase bench
seat divider mouse wall windowsill handicap bar cushion hair dryer ceiling
storage bin closet wall scale plant bulletin board purse table mini fridge piano
coffee maker bathroom stall door power outlet coffee table fireplace vent nightstand dumbbell closet
dishwasher stair rail decoration stairs mini fridge shower floor bathroom vanity oven cabinet
bar bathroom cabinet sign armchair water cooler water pitcher laptop luggage cup
toaster closet rod projector cabinet shower door bowl shower wall bar laundry hamper
ironing board structure vacuum cleaner bathroom vanity pillar paper bag desk pipe light switch
fireplace coat rack candle bathroom stall ledge alarm clock computer tower bathroom stall cd case
kitchen counter storage organizer plunger mirror furniture music stand soap dispenser blinds backpack
toilet paper dispenser stuffed animal blackboard cart laundry detergent container toilet paper dispenser windowsill
mini fridge headphones trash can decoration dumbbell bicycle coffee table box
tray broom stair rail closet door tube light dish rack book
toaster oven guitar case box vacuum cleaner cd case clothes guitar mailbox
toilet seat cover dispenser hair dryer towel dish rack closet rod machine seat sofa chair
furniture water bottle door range hood coffee kettle furniture clock shower curtain
cart purse clothes projector screen shower head stair rail alarm clock bulletin board
storage container vent whiteboard divider keyboard piano toilet paper holder board crate
tissue box water pitcher bed bathroom counter case of water bottles floor file cabinet tube
crate bowl floor laundry hamper coat rack bucket ceiling light window
dish rack paper bag bathtub bathroom stall door folded chair stool ladder power outlet
range hood alarm clock desk ceiling light fire alarm door paper towel dispenser power strip
dustpan laundry detergent wardrobe trash bin power strip sign shower floor bathtub
handicap bar object clothes dryer bathroom cabinet calendar recycling bin stuffed animal column
mailbox ceiling light radiator structure poster shower water cooler fire alarm
music stand dumbbell shelf storage organizer luggage jacket coffee kettle storage container
bathroom counter tube potted plant bottle kitchen counter
bathroom vanity cd case mattress
laundry hamper coffee kettle
trash bin shower head
keyboard piano case of water bottles
folded chair fire alarm
luggage power strip
mattress calendar
poster
potted plant
Appendix Appendix B Incremental Scenarios Phases
Table 7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset. The three scenarios, Split_A, Split_B, and Split_C, are each divided into three tasks: Task 1, Task 2, and Task 3. Notably, the order of classes in these tasks is random.
Appendix Appendix C Qualitative Results
In this section, we present a qualitative comparison of the proposed framework with the baseline method. Figure 4 illustrates the results on the Split_A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach. As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline. For example, in row 1, the baseline method fails to segment the sink, while in row 3, the sofa instance is missed. Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_A  evaluation after learning all tasks.
In Figure 5, we present the results on Split_B, highlighting instances where the baseline method underperforms, marked with red circles. For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances. Similarly, in row 5, the washing machine is segmented into two instances by the baseline. In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_B  evaluation after learning all tasks.
Similarly, Figure 6 highlights the results on Split_C, where classes are encountered in random order. The comparison emphasizes the advantages of our method, as highlighted by red circles. The baseline method often misses instances or splits a single instance into multiple parts. In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_C  evaluation after learning all tasks.