### v2

# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

Abstract
While 3D instance segmentation (3DIS) has advanced significantly, existing methods typically assume that all object classes are known in advance and are uniformly distributed. However, this assumption is unrealistic in dynamic, real-world environments where new classes emerge gradually and exhibit natural imbalance. Although some approaches have addressed class emergence, they often overlook class imbalance, resulting in suboptimal performance – particularly on rare categories. To tackle this challenge, we propose CLIMB-3D, a unified framework for CLass-incremental Imbalance-aware 3DIS. Building upon established exemplar replay (ER) strategies, we show that ER alone is insufficient to achieve robust performance under constrained memory conditions. To mitigate this, we introduce a novel pseudo-label generator (PLG) that extends supervision to previously learned categories by leveraging predictions from a frozen prior model. Despite its promise, PLG tends to bias towards frequent classes. Therefore, we propose a class-balanced re-weighting (CBR) scheme, that estimates object frequencies from pseudo-labels and dynamically adjusts training bias – without requiring access to past data. We design and evaluate three incremental scenarios for 3DIS on the challenging ScanNet200 dataset, and additionally on semantic segmentation on ScanNetV2. Our approach achieves state-of-the-art results, surpassing prior work by up to 16.76% mAP for instance segmentation and approximately 30% mIoU for semantic segmentation, demonstrating strong generalization across both frequent and rare classes.
1 Introduction
Figure 1: Overview of the CI-3DIS setting: new classes are introduced incrementally as tasks progress. After each task, the model must recognize both previously learned and newly added classes. For instance, at Task 2, classes like Pillow, Coffee Table, and Sofa Chair are introduced, while the model retains the ability to detect earlier classes such as Table, Chair, and Couch.
3D instance segmentation (3DIS) is an essential task in computer vision with widespread applications in graphics, robotics, and augmented reality [boudjoghra20243d, lai2022stratified]. It involves identifying and segmenting individual objects in 3D space, enabling precise object boundaries and class labels that enhance scene understanding and perception in dynamic environments.
In recent years, a variety of methods for 3DIS have been proposed, including top-down [wang2018sgpn, jia2021scaling, zhang2021point], bottom-up [hou20193dsis, yang2019learning], and transformer-based approaches [Schult23ICRA]. These methods have demonstrated strong performance under the traditional setting, which assumes that all object classes are available and well-balanced during training. However, these assumptions limit their applicability in real world settings, where new categories gradually emerge over time, often exhibiting natural class imbalance.
These limitations in current 3DIS methods underscore the need for a Class-Incremental Learning (CIL) [rebuffi2017icarl, de2021continual] framework capable of learning new categories as they appear, while maintaining accuracy on previously learned classes – especially rare ones, which are more susceptible to without catastrophic forgetting [mccloskey1989catastrophic]. Unfortunately, most existing research in CIL has focused on the 2D domain for image classification [rebuffi2017icarl, li2017learning, aljundi2018memory, serra2018overcoming], with limited extensions to object detection [liu2023continual, joseph2021towards, shmelkov2017incremental] and semantic segmentation [cermelli2020modeling, cermelli2022incremental, douillard2021plop]. Extensions to 3D point clouds remain limited, with most work focusing on object-level classification rather than scene-level segmentation [dong2021i3dol, liu2021l3doc, chowdhury2021learning]. Recent works on scene-level class-incremental 3DIS (CI-3DIS) [boudjoghra20243d] and class-incremental semantic segmentation (CI-3DSS) [Yang_2023_CVPR] show promise, but typically rely on large exemplar buffers [rolnick2019experience] Moreover, they typically overlook class imbalance, limiting real-world applicability.
To address this, we propose CLIMB-3D, a unified framework for CLass-incremental IMBalance-aware 3DIS that jointly tackles catastrophic forgetting and class imbalance. The framework begins with Exemplar Replay (ER), storing a small number of samples from past classes for later replay. However, this alone does not yield promising results under the strict memory constraints needed CI-3DIS. To address this, we introduce a Pseudo-Label Generator (PLG), which uses a frozen previous task model to generate supervision for earlier classes. However, we observed that, PLG tends to favour frequent classes while ignoring under-represented ones. To mitigate this, we propose a Class-Balanced Re-weighting (CBR), which estimates class frequencies from the pseudo-labels to enhance the learning of rare classes without accessing past data. Together, these components form a practical and effective solution for real-world 3DIS.
To evaluate CLIMB-3D in a realistic incremental setup, we design three benchmark scenarios using the ScanNet200 [rozenberszki2022language] dataset, which contains 200 classes with a natural long-tail distribution. These scenarios simulate real-world conditions where new classes gradually emerge, following inherently imbalanced distributions. Specifically, new classes appear in groups based on one of three criteria: (A) object frequency, (B) semantic similarity, or (C) random grouping without a specific pattern. Our experiments on these scenarios for incremental 3DIS show that CLIMB-3D effectively mitigates forgetting of previous tasks while improving performance across both frequent and rare classes, delivering consistent gains over prior methods.
In summary, our contributions are threefold: (i) We introduce a novel problem setting of imbalanced class-incremental 3DIS, along with a simple and effective method to mitigate catastrophic forgetting and ensure balanced learning; (ii) We establish three benchmark scenarios that realistically model continuous object emergence with natural class imbalance; and (iii) We validate our framework’s effectiveness, achieving SOTA results with improvements up to 16.76% mAP over existing baselines [^1] Code will be available at: https://github.com/vgthengane/CLIMB3D .
2 Related Work
This section reviews the current literature on 3D instance segmentation and incremental learning methods, including the limited work addressing incremental learning for 3D scene-level tasks.
2.1 3D Instance Segmentation
Various approaches have been proposed for 3D instance segmentation. One common approach adopts a bottom-up pipeline, in which an embedding in the latent space is learned to facilitate the clustering of object points [wang2018sgpn, lahoud20193d, jiang2020pointgroup, zhang2021point, chen2021hierarchical, han2020occuseg, he2021dyco3d, liang2021instance]. These methods are also known as grouping-based or clustering-based methods. Other methods use a top-down approach, also known as proposal-based methods, where 3D bounding boxes are first detected, then the object region is segmented within the box [yang2019learning, hou20193dsis, liu2020learning, yi2019gspn, engelmann20203d]. Recently, the transformer architecture [vaswani2017attention] has also been used for the task of 3D instance segmentation [Schult23ICRA, sun2022superpoint], motivated by work in 2D [cheng2022masked, cheng2021per]. While these methods propose various models for improving the quality of the object segments, they rely on the availability of annotations for all object categories. On the other hand, we target learning in a progressive manner, in which new semantic annotation is provided and past data is inaccessible.
In order to reduce the annotation cost for 3D instance segmentation, various methods propose weakly supervised alternatives to methods that use dense annotations [xie2020pointcontrast, hou2021exploring, chibane2022box2mask]. While these methods improve the ability to learn from a small set of annotated examples, they rely on a fixed set of semantic labels, so they are prone to catastrophic forgetting in an incremental setting.
2.2 Incremental Learning
Incremental, lifelong, or continual learning methods aim to train a machine learning model sequentially to avoid “catastrophic forgetting” which is caused by training the model on a set of data and later training on another set of data. There are several methods have been proposed for this paradigm, these methods can be divided into three categories: (i) Model Regularization [kirkpatrick2017overcoming, aljundi2018memory, li2017learning] methods limit the plasticity of model parameters to avoid catastrophic forgetting of previous tasks. These methods include weight regularization such as EWC [serra2018overcoming] and function regularization such as knowledge distillation [hinton2015distilling]. (ii) Exemplar replay approaches either create a subset of the past task data or generate samples using generative models to avoid privacy concerns and save those in memory to replay while learning new tasks. This method is effective in more challenging settings and datasets [rebuffi2017icarl, kamra2017deep, buzzega2020dark, cha2021co2l]. (iii) Dynamic network expansion-based method learns a new task by either dynamically expanding the model [rusu2016progressive, li2019learn, zhao2022deep] or by creating a subset of the model [zhao2022deep, wang2020learn, ke2020continual, rajasegaran2019random] to learn to cater for a new task.
Recent approaches to 3D class-incremental segmentation, such as [Yang_2023_CVPR] and [su2024balanced], have made some initial contributions. However, these methods often fall short in performance as they do not leverage state-of-the-art 3D segmentation models and are primarily focused on semantic segmentation, while our work emphasizes object-level instance segmentation. Kontogianni et al. [kontogianni2024continual] propose a general online continual learning framework and evaluate it on 3D dataset segmentation. Similarly, [boudjoghra20243d] addresses the open-world 3D incremental learning problem but relies heavily on an extensive memory buffer. In contrast, our work introduces a dedicated continual learning framework for 3D instance segmentation, with a focus on effective knowledge transfer from previous tasks, while also accounting for the challenges posed by infrequent class occurrences.
3 Methodology
Figure 2: Overview of CLIMB-3D for CI-3DIS. The model incrementally learns new classes in sequential phases using ER to mitigate forgetting, a PLG for supervision on old classes, and CBR to handle class imbalance.
3.1 Problem Formulation
3DIS. The objective of this task is to detect and segment each object instance within a given point cloud. Formally, the training dataset is denoted by $\mathcal{D}=\{(\mathbf{P}_{i},\mathbf{Y}_{i})\}_{i=1}^{N}$, where $N$ is the number of training samples. Each input $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ is a colored point cloud consisting of $M$ points, where each point is represented by its 3D coordinates and RGB values. The corresponding annotation $\mathbf{Y}_{i}=\{(m_{i,j},c_{i,j})\}_{j=1}^{J_{i}}$ contains $J_{i}$ object instances, where $m_{i,j}\in\{0,1\}^{M}$ is a binary mask indicating the points belonging to the $j$-th instance, and $c_{i,j}\in\mathcal{C}=\{1,\dots,C\}$ is the semantic class label of that instance. Given $\mathbf{P}_{i}$, the model $\Phi$ predicts instance-level outputs $\hat{\mathbf{Y}}_{i}=\{(\hat{m}_{i,j},\hat{c}_{i,j})\}_{j=1}^{\hat{J}_{i}}$, where $\hat{m}_{i,j}$ and $\hat{c}_{i,j}$ denote the predicted mask and category label for the $j$-th instance, respectively. The number of predicted instances $\hat{J}_{i}$ varies depending on the model’s inference. The model is optimized using below objective:
$\mathcal{L}_{\text{3DIS}}(\mathcal{D};\Phi)$ $=\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{P},\mathbf{Y})\in\mathcal{D}}\frac{1}{|\mathbf{Y}|}\sum_{(m_{j},c_{j})\in\mathbf{Y}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j},\hat{m}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j},\hat{c}_{j})\right)$ (1)
where $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{mask}}$ are the average classification and mask losses over all instances, and $\lambda$ controls their trade-off.
CI-3DIS. Unlike conventional 3DIS, CI-3DIS involves sequential learning, where object categories are introduced incrementally over $T$ training tasks. Each task $t$ introduces a disjoint set of classes $\mathcal{C}^{t}$, with $\mathcal{C}=\bigcup_{t=1}^{T}\mathcal{C}^{t}$ and $\mathcal{C}^{t}\cap\mathcal{C}^{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. At each task $t$, the model receives a dataset $\mathcal{D}^{t}=\{(\mathbf{P}_{i},\mathbf{Y}_{i}^{t})\}_{i=1}^{N}$, where each coloured point cloud $\mathbf{P}_{i}\in\mathbb{R}^{M\times 6}$ contains $M$ points, and the corresponding annotation $\mathbf{Y}_{i}^{t}=\{(m_{i,j}^{t},c_{i,j}^{t})\}_{j=1}^{J_{i}^{t}}$ includes instance masks $m_{i,j}^{t}\in\{0,1\}^{M}$ and semantic labels $c_{i,j}^{t}\in\mathcal{C}^{t}$. The model $\Phi^{t}$, initialized from $\Phi^{t-1}$, is trained on $\mathcal{D}^{t}$ to predict instance-level outputs $\hat{\mathbf{Y}}_{i}^{t}=\{(\hat{m}_{i,j}^{t},\hat{c}_{i,j}^{t})\}_{j=1}^{\hat{J}_{i}^{t}}$, where each predicted class label $\hat{c}_{i,j}^{t}$ belongs to the cumulative set of all classes observed so far: $\bigcup_{k=1}^{t}\mathcal{C}^{k}$. The optimisation objective for the CI-3DIS setting is defined as:
$\mathcal{L}_{\text{CI-3DIS}}(\mathcal{D}^{t};\Phi^{t})$ $=\frac{1}{|\mathcal{D}^{t}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in\mathcal{D}^{t}}\frac{1}{|\mathbf{Y}|}\sum_{(m^{t}_{j},c^{t}_{j})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m^{t}_{j},\hat{m}^{t}_{j})+\lambda\cdot\mathcal{L}_{\text{cls}}(c^{t}_{j},\hat{c}^{t}_{j})\right).$ (2)
The key challenge in this setup is to learn new classes without forgetting those encountered in previous tasks, despite limited supervision and the absence of past labels.
3.2 Method: CLIMB-3D
Overview. As illustrated in Fig. 2 and introduced in Sec. 3.1, the proposed framework for CI-3DIS follows a phase-wise training strategy. In each phase, the model is exposed to a carefully selected subset of the dataset, designed to simulate real-world scenarios discussed in Sec. 3.3. Naïvely training on such phased data leads to catastrophic forgetting [mccloskey1989catastrophic], where the model forgets previously acquired knowledge when learning new tasks. To address this, we first incorporate Exemplar Replay (ER) [buzzega2020dark], inspired by 2D [rebuffi2017icarl, li2017learning, aljundi2018memory] and 3D [boudjoghra20243d] incremental setting, to mitigate forgetting by storing a small set of representative samples from earlier phases. However, ER alone is insufficient for achieving robust performance. We therefore introduce two additional components: a Pseudo-Label Generator (PLG), which leverages a frozen model to generate supervision signals for previously seen classes, and a Class-Balanced Re-weighting (CBR) module that compensates for class imbalance across phases. Each of these components is described in detail below  [^2] For clarity, modifications introduced by each component – ER, PLG, and CBR – are highlighted in green, blue, and red, respectively..
Exemplar Replay (ER). Inspired by Buzzega et al. [buzzega2020dark], ER addresses the issue of catastrophic forgetting by allowing the model to retain a small portion of data from earlier phases. At phase $t$, the model is trained not only on the current task data $\mathcal{D}^{t}$ but also on a set of exemplars $\mathcal{E}^{1:t-1}$ collected from previous phases. These exemplars are accumulated incrementally, where $\mathcal{E}^{1:t-1}=\mathcal{E}^{1}\cup\dots\cup\mathcal{E}^{t-1},$ and $\mathcal{E}^{t}$ denotes the exemplar subset stored after phase $t$, with $|\mathcal{E}^{t}|$ as the exemplar replay size. We define the combined replay dataset as: ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}=\mathcal{D}^{t}\cup\mathcal{E}^{1:t-1}.$ Epoch training proceeds in two stages: first on the current data $\mathcal{D}^{t}$, followed by replay using $\mathcal{E}^{1:t-1}$. The updated objective function for CI-3DIS, based on LABEL:eq:ci3dis-loss, becomes:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|\mathbf{Y}^{t}|}\sum_{(m_{j}^{t},c_{j}^{t})\in\mathbf{Y}^{t}}$
$\quad\left(\mathcal{L}_{\text{mask}}(m_{j}^{t},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}(c_{j}^{t},\hat{c}_{j}^{t})\right).$ (3)
While previous CI-3DIS approaches use large exemplar sets to mitigate forgetting [boudjoghra20243d], such strategies become impractical under memory constraints. Consequently, although exemplar replay (ER) aids knowledge retention, it alone proves insufficient for robust performance with a limited exemplar budget. We therefore propose PLG and CBM to better preserve past representations and enhance generalisation across phases.
Pseudo-Label Generator (PLG). In the CI-3DIS setting, at phase $t$ (for $t>1$), although we no longer have access to the ground-truth labels for previously seen classes, the model from the previous phase, $\Phi^{t-1}$, remains available and retains knowledge acquired from past tasks. We use this model to generate pseudo-labels for previous classes, thereby providing approximate past supervision during training. This allows the current model $\Phi^{t}$ to retain past knowledge and reduces forgetting, even without access to original annotations.
Given a point cloud and label pair $(\mathbf{P},\mathbf{Y}^{t})$ from $\mathcal{D}^{t}$, the previous model model $\Phi^{t-1}$ generates pseudo-labels for the previously learned classes as $\hat{\mathbf{Y}}^{1:t-1}=\Phi^{t-1}(\mathbf{P})$, which we denote as $\tilde{\mathbf{Y}}^{t}$ for brevity. However, we observe that using all predictions from $\Phi^{t-1}$ introduces noisy or incorrect labels, which degrade overall performance. To mitigate this, we select the top-$K$ most confident instance predictions, denoted as $\tilde{\mathbf{Y}}^{t}_{K}$, and concatenate them with the ground-truth labels of the current task to obtain the supervision set: ${\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}$. At the end, Using this supervision set, the model is optimized via the following updated objective:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}})\in{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{\mathbf{Y}}^{t}}}$
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{m}_{j}^{t}},\hat{m}_{j}^{t})+\lambda\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\bar{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (4)
where the loss is computed over both the current task’s ground-truth labels and the top-$K$ confident pseudo-labels from previous phases, enabling the model to retain prior knowledge while adapting to new classes.
Class-Balanced Re-weighting (CBR). While ER and PLG contribute to preserving prior knowledge, we observe that the model’s predictions are biased toward frequently occurring classes, leading to forgetting of rare categories. This issue is further reinforced by the top-$K$ pseudo-labels selected by PLG, which predominantly represent dominant classes in the data. To alleviate this issue, we introduce a Class-Balanced Re-weighting (CBR) scheme that adjusts for class imbalance based on frequency of object occurrence. However, at task $t$, we only have access to the dataset $\mathcal{D}^{t}$ and its corresponding label distribution $p^{t}(c)$ over classes $c\in\mathcal{C}^{t}$. The datasets and class distributions from previous tasks $\{\mathcal{D}^{i},p^{i}(c)\}_{i=1}^{t-1}$ are no longer accessible – making it challenging to directly account for class imbalance across all seen tasks. Therefore, we propose to leverage the pseudo-label predictions of the frozen model $\Phi^{t-1}$ as a proxy for the class distribution across previously learned categories.
During each training iteration, we apply the previous model $\Phi^{t-1}$ to the current input point cloud $\mathbf{P}$ to generate pseudo-labels $\tilde{\mathbf{Y}}^{t}$ , from which we accumulate class-wise frequency statistics corresponding to prior tasks. Let $\mathcal{C}^{1:t}=\bigcup_{i=1}^{t}\mathcal{C}^{i}$ represent the union of all classes seen up to task $t$. At the end of each epoch, we compute the overall class frequency $\mathbf{f}\in\mathbb{R}^{|\mathcal{C}^{1:t}|}$ by combining the pseudo-label distribution $\tilde{p}^{t}(c)$ derived from $\tilde{\mathbf{Y}}^{t}$ with the observed ground-truth label distribution $p^{t}(c)$ from the current task. The resulting frequency vector $\mathbf{f}=[f(c)]_{c\in\mathcal{C}^{1:t}}$ is then used to compute class-wise $\mathbf{w}=[w(c)]_{c\in\mathcal{C}^{1:t}}$ weights for balanced pseudo-label selection and to tune current model to favour rare classes.
During the next epoch, the model $\Phi^{t-1}$ produces soft pseudo-label predictions $\tilde{\mathbf{Y}}^{t}$ , where each component $\tilde{\mathbf{Y}}^{t}[c]$ denotes the confidence score for class $c\in\mathcal{C}^{1:t-1}$. To promote balanced pseudo-label selection, we apply class-wise re-weighting using $w(c)=\frac{1}{f(c)+\epsilon}$, where $\epsilon$ is a small constant for numerical stability. The adjusted predictions are computed by applying the class-wise weights as: $\tilde{\mathbf{Y}}^{t,w}[c]=w(c)\cdot\tilde{\mathbf{Y}}^{t}[c],\quad\forall c\in\mathcal{C}^{1:t-1}.$ We then perform top-$K$ selection over both the original and the re-weighted scores, yielding two pseudo-label sets: $\tilde{\mathbf{Y}}^{t}_{K}$ and $\tilde{\mathbf{Y}}^{t,w}_{K}$, respectively. The final supervision label set $\overline{\mathbf{Y}}^{t}$ is constructed by concatenating the ground-truth labels $\mathbf{Y}^{t}$ with both sets of pseudo-labels: ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}=\mathbf{Y}^{t}\mathbin{\|}{\color[rgb]{0,0.19921875,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.19921875,0.703125}\tilde{\mathbf{Y}}^{t}_{K}}\mathbin{\|}{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\tilde{\mathbf{Y}}^{t,w}_{K}}.$
Although the above re-weighting mitigates bias from prior tasks, it does not address class imbalance within the current task $t$. To resolve this, we extend the class-balanced weights $w_{c}$ to incorporate statistics from both previously seen and current classes, i.e., over $\mathcal{C}^{1:t}$ rather than only $\mathcal{C}^{1:t-1}$, and denote the updated weights as $w_{c}^{\prime}$, which are then integrated into the training objective. After incorporating all components, the modified objective in Eq. 1 becomes the final updated objective used in CLIMB-3D:
$\mathcal{L}_{\text{CI-3DIS}}({\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}};\Phi^{t})$ $=\frac{1}{|{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}|}\sum_{(\mathbf{P},\mathbf{Y}^{t})\in{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\mathcal{D}^{t}_{\text{ER}}}}\frac{1}{|{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}|}\sum_{({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}})\in{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{\mathbf{Y}}^{t}}}$
$\quad\left(\mathcal{L}_{\text{mask}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{m}_{j}^{t}},\hat{m}_{j}^{t})+{\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}\cdot\mathcal{L}_{\text{cls}}({\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\overline{c}_{j}^{t}},\hat{c}_{j}^{t})\right),$ (5)
where ${\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime\prime}_{c}}={\color[rgb]{0.62890625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0,0}\mathbf{w}^{\prime}_{c}}\cdot\lambda_{\mathrm{cls}}$ denotes the scaled weight vector, and $\lambda_{\mathrm{cls}}$ is a loss-balancing hyperparameter. To address the class-imbalance across tasks, our CBR module jointly re-weights pseudo-label selection and classification loss, ensuring balanced supervision across all seen categories, including rare ones. This unified strategy reduces bias and enhances performance, particularly for under-represented classes.
3.3 Benchmarking Incremental Scenarios
While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution – an assumption that rarely holds in real-world settings. In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size. To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity. The design of these scenarios is detailed below; for more information and illustration, please refer Appendix Appendix C of the supplementary material.
Split_A: Frequency Scenarios. This scenario acknowledges that datasets are often labelled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rare categories.
Split_B: Semantic Scenarios. Beyond frequency, semantic similarity is crucial in real-world deployments. While objects often share visual or functional traits, models may encounter semantically different categories in new environment. To simulate this challenge, we introduce the Split_Bscenario. Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time. This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones. Unlike the Split_Ascenario, which organises learning based on category frequency, the Split_Bscenario may contain both frequent and infrequent categories within a single task, focusing instead on semantic continuity and transfer.
Split_C: Random Scenarios. In some cases, data labelling is driven by the availability of objects rather than predefined criteria. To capture this, we introduce the Split_Cscenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance. By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
By designing these incremental scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
4 Experiments
Table 1: Comparison between the proposed method and the baseline in the 3DIS setting, evaluated using mAP25, mAP50, mAP, and FPP after training across all phase.
Scenarios Methods Average Precision $\uparrow$ FPP $\downarrow$
mAP25 mAP50 mAP mAP25 mAP50
Baseline 16.46 14.29 10.44 51.30 46.82
Split_A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
Baseline 17.22 15.07 10.93 46.27 42.1
Split_B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
Baseline 25.65 21.08 14.85 31.68 28.84
Split_C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
Table 2: Comparison between the proposed method and existing baselines in the 3DSS setting on ScanNetV2 [dai2017scannet], evaluated using mIoU.
Methods Phase=1 Phase=2 All
EWC [serra2018overcoming] 17.75 13.22 16.62
LwF [li2017learning] 30.38 13.37 26.13
Yang Yang_2023_CVPR 34.16 13.43 28.98
CLIMB-3D (Ours) 69.39 32.56 59.38
Table 3: Ablation study showing the impact of each component ER, PLG, and CBR in a three-phase setup. Each split (s) and number corresponds to data introduced at that phase. The All column reports performance over all classes seen so far. Joint training (Oracle) is highlighted in gray, and the full model in orange. Best results are in bold.
Row Modules p=1 $\uparrow$ p=2 $\uparrow$ p=3 $\uparrow$ FPP $\downarrow$
s1 s1 s2 All s1 s2 s3 All
1. Oracle - - - - 55.14 30.77 25.30 37.68 -
2. Naïve 56.82 0.00 28.09 14.15 0.00 0.00 19.67 5.80 56.82
3. + ER 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28 46.44
4. + PLG 56.82 50.00 34.39 42.13 49.78 11.41 26.47 29.28 7.04
5. + CBR 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05 2.63
4.1 Setup
Datasets. We evaluate our method on the ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance – making it ideal for simulating and assessing real-world scenarios. In addition, we benchmark our approach against existing incremental learning methods using the ScanNetV2 [dai2017scannet] in a 3DSS setting. For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR].
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3DIS that provides a comprehensive measure of segmentation quality by accounting for both precision and recall. For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments. To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual]. FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
Incremental Scenarios. As discussed in Section 3.3, we design three incremental scenarios – Split_A, Split_B, and Split_C– each with three tasks, based on object frequency, semantic similarity, and random grouping, respectively. For Split_A, we follow the class split provided by rozenberszki2022language, with a class distribution of 66-68-66 across the three tasks. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, classes are randomly shuffled and split into three sets, with 67-67-66 categories per split. These scenarios comprehensively evaluate our approach’s performance and generalization in varied real-world conditions.
Implementation Details. For all the experiments we use Mask3D [Schult23ICRA] with the Minkowski Res16UNet34C [choy20194d] as a feature backbone. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA], with using the AdamW optimizer [loshchilov2017decoupled] and a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, learning is divided into three phases, introducing one split per phase across the three scenarios described in Sec. 3.3. For each phase, we train for 500, 300, and 300 epochs, respectively, except for the Oracle setting (Row 1 in Tab. 3), which is trained once for 600 epochs. We retain the same hyperparameters throughout, modifying only the number of training epochs. The exemplar replay size is fixed at $|\mathcal{E}^{t}|=50$ scenes.
4.2 Results and Discussion
To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task. In contrast, for the 3DSS setting, where prior incremental baselines do exist, we compare our method against three existing approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
Results on CI-3DIS. Tab. 1 compares CLIMB-3D against the ER baseline across three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP. In the Split_Ascenario, characterized by significant distribution shifts, it achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, while drastically reducing forgetting with FPP scores of 3.44% and 2.63%, over 47 points lower than the baseline. Under the Split_Bscenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%. For the Split_Cscenario, featuring increasing geometric complexity, it maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points. Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, validating the method’s effectiveness for scalable continual 3D semantic understanding.
Results on CI-3DSS. Although our primary focus is on CI-3DIS, we additionally evaluate our method under the CI-3DSS setting to enable comparison with existing approaches. For this, we adapt our predictions by assigning each point the label corresponding to the highest-confidence mask and exclude background classes (floor and wall), as these are not part of the semantic segmentation targets. Tab. 2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall. Despite being designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains: +35.23% mIoU in Phase 1 and +19.1% in Phase 2. Overall, it achieves 59.38% mIoU – significantly higher than the $\sim$30mIoU of prior methods – highlighting its robustness and transferability.
For a detailed per-phase analysis, evaluation of rare classes, and qualitative result comparisons, please refer to Appendix Appendix A, Appendix Appendix B, and Appendix Appendix D, respectively, in the supplementary material.
4.3 Ablation
We conduct an ablation study to evaluate the contribution of each component in our framework. As an upper bound, we report the performance of an Oracle model trained jointly on the full dataset. For the incremental setting, we follow the previously defined splits and first train the model naively across phases. We then incrementally add each module to isolate its impact. Tab. 3 presents results on the Split_Ascenario using both $\text{mAP}_{50}$  and FPP metrics.
Naïve Training. When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2. The model completely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and a high FPP of 56.82.
Effect of ER. Adding exemplar replay (row 3) partially alleviates forgetting by maintaining a buffer of past examples. This yields notable gains for s1 in phase 2 (+18.51%) and phase 3 (+10.38%), and also improves s2 in phase 3 (+9.43%). However, the overall forgetting remains substantial (FPP: 46.44), showing ER alone is insufficient.
Effect of PLG. The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from earlier phase, facilitates knowledge retention and forward transfer. As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks. For s1, it improves $\text{mAP}_{50}$ by 31.49% in phase 2 and 39.40% in phase 3 over exemplar replay. Overall, PLG yields a 15.00% increase in performance and reduces forgetting by 39.40%.
Effect of CBR. Finally, class-balanced re-weighting (CBR) mitigates class imbalance both during pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency. As shown in row 5, CBR enhances s1 retention over PLG in phases 2 and 3 (+4.67% and +4.41%) and further improves performance on s2 and s3. It also achieves the lowest FPP of 2.63, reflecting effective forgetting mitigation. Overall, CBR offers the best trade-off between retaining past knowledge and learning new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
5 Conclusion
We address the critical challenge of catastrophic forgetting in class-incremental 3D instance segmentation by proposing a modular framework that combines exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR). Our method incrementally adapts to new classes while preserving knowledge from previous tasks without requiring access to the full dataset. Through extensive experiments on a three-phase benchmark, we demonstrate the individual and combined effectiveness of each component. ER provides a easy knowledge storage, PLG enhances knowledge retention via pseudo-label supervision, and CBR further refines learning by mitigating class imbalance. Together, these modules significantly reduce forgetting and improve overall segmentation performance across all phases. By balancing stability and plasticity, our method contributes to continual 3D scene understanding and establishing a new baseline for future research.
Supplementary Material
Table 4: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios. Each scenario is trained in three phases ($\mathtt{phase}=1,2,3$) by introducing a single split s at a time. The results highlighted in orange are with the proposed method, and the best results for each scenario are in bold.
Scenarios Methods phase=1 phase=2 phase=3
s1 s1 s2 All s1 s2 s3 All
Baseline 56.82 18.51 32.81 25.72 10.38 9.43 24.27 14.28
Split_A CLIMB-3D 56.82 54.67 33.75 44.13 54.19 12.02 26.55 31.05
Baseline 51.57 13.32 42.21 24.53 9.55 12.45 26.78 15.07
Split_B CLIMB-3D 51.57 46.74 37.45 43.13 46.06 15.95 26.68 31.56
Baseline 36.40 7.74 37.62 22.32 7.55 15.96 40.41 21.08
Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
paper towel dispenser 2 73.10 74.90
recycling bin 3 55.80 60.50
ladder 5 53.90 57.10
trash bin 7 31.50 57.30
bulletin board 8 23.30 38.20
shelf 11 48.00 50.50
dresser 12 44.00 55.80
copier 12 93.30 94.50
object 12 3.10 3.30
stairs 13 51.70 67.70
bathtub 16 80.30 86.60
oven 16 1.50 3.30
divider 18 36.40 45.00
column 20 57.30 75.00
Average - 46.66 54.98
Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
piano 1 7.10 59.40
bucket 1 21.10 31.50
laundry basket 1 3.80 17.40
dresser 2 55.00 55.40
paper towel dispenser 2 32.50 35.50
cup 2 24.70 30.30
bar 2 35.40 39.50
divider 2 28.60 42.40
case of water bottles 2 0.00 1.70
shower 3 0.00 45.50
mirror 8 56.00 68.80
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
bathroom counter 3 3.90 20.30
ottoman 4 32.60 36.20
storage bin 3 5.10 10.50
dishwasher 3 47.40 66.20
trash bin 4 1.10 2.70
backpack 5 74.50 76.70
copier 5 94.00 96.80
sofa chair 6 14.10 43.50
file cabinet 6 49.20 57.60
tv stand 7 67.70 68.60
mirror 8 56.00 68.80
blackboard 8 57.10 82.80
clothes dryer 9 1.70 3.20
toaster 9 0.10 25.90
wardrobe 10 22.80 58.80
jacket 10 1.20 4.10
Average - 32.08 44.21
Figure 3: Tasks are grouped into incremental scenarios based on object frequency, semantic similarity, and random assignment.  ,  , and   denote different tasks; shapes indicate object categories;   marks the background. Left: Grouped by category frequency. Middle: Grouped by semantic similarity (e.g., similar shapes). Right: Randomly grouped, mixing semantic and frequency variations.
Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels. The three scenarios Split_A, Split_A, and Split_C  are each divided into three tasks: Task 1, Task 2, and Task 3.
Split_A Split_B Split_C
Task 1 Task 2 Task 3 Task 1 Task 2 Task 3 Task 1 Task 2 Task 3
chair wall pillow tv stand cushion paper broom fan rack
table floor picture curtain end table plate towel stove music stand
couch door book blinds dining table soap dispenser fireplace tv bed
desk cabinet box shower curtain keyboard bucket blanket dustpan soap dish
office chair shelf lamp bookshelf bag clock dining table sink closet door
bed window towel tv toilet paper guitar shelf toaster basket
sink bookshelf clothes kitchen cabinet printer toilet paper holder rail doorframe chair
toilet curtain cushion pillow blanket speaker bathroom counter wall toilet paper
monitor kitchen cabinet plant lamp microwave cup plunger mattress ball
armchair counter bag dresser shoe paper towel roll bin stand monitor
coffee table ceiling backpack monitor computer tower bar armchair copier bathroom cabinet
refrigerator whiteboard toilet paper object bottle toaster trash bin ironing board shoe
tv shower curtain blanket ceiling bin ironing board dishwasher radiator blackboard
nightstand closet shoe board ottoman soap dish lamp keyboard vent
dresser computer tower bottle stove bench toilet paper dispenser projector toaster oven bag
stool board basket closet wall basket fire extinguisher potted plant paper bag paper
bathtub mirror fan couch fan ball coat rack structure projector screen
end table shower paper office chair laptop hat end table picture pillar
dining table blinds person kitchen counter person shower curtain rod tissue box purse range hood
keyboard rack plate shower paper towel dispenser paper cutter stairs tray coffee maker
printer blackboard container closet oven tray fire extinguisher couch handicap bar
tv stand rail soap dispenser doorframe rack toaster oven case of water bottles telephone pillow
trash can radiator telephone sofa chair piano mouse water bottle shower curtain rod decoration
stairs wardrobe bucket mailbox suitcase toilet seat cover dispenser ledge trash can printer
microwave column clock nightstand rail storage container shower head closet wall object
stove ladder stand washing machine container scale guitar case cart mirror
bin bathroom stall light picture telephone tissue box kitchen cabinet hat ottoman
ottoman shower wall pipe book stand light switch poster paper cutter water pitcher
bench mat guitar sink light crate candle storage organizer refrigerator
washing machine windowsill toilet paper holder recycling bin laundry basket power outlet bowl vacuum cleaner divider
copier bulletin board speaker table pipe sign plate mouse toilet
sofa chair doorframe bicycle backpack seat projector person paper towel roll washing machine
file cabinet shower curtain rod cup shower wall column candle storage bin laundry detergent mat
laptop paper cutter jacket toilet bicycle plunger microwave calendar scale
paper towel dispenser shower door paper towel roll copier ladder stuffed animal office chair wardrobe dresser
oven pillar machine counter jacket headphones clothes dryer whiteboard bookshelf
piano ledge soap dish stool storage bin broom headphones laundry basket tv stand
suitcase light switch fire extinguisher refrigerator coffee maker guitar case toilet seat cover dispenser shower door closet rod
recycling bin closet door ball window dishwasher dustpan bathroom stall door curtain plant
laundry basket shower floor hat file cabinet machine hair dryer speaker folded chair counter
clothes dryer projector screen water cooler chair mat water bottle keyboard piano suitcase bench
seat divider mouse wall windowsill handicap bar cushion hair dryer ceiling
storage bin closet wall scale plant bulletin board purse table mini fridge piano
coffee maker bathroom stall door power outlet coffee table fireplace vent nightstand dumbbell closet
dishwasher stair rail decoration stairs mini fridge shower floor bathroom vanity oven cabinet
bar bathroom cabinet sign armchair water cooler water pitcher laptop luggage cup
toaster closet rod projector cabinet shower door bowl shower wall bar laundry hamper
ironing board structure vacuum cleaner bathroom vanity pillar paper bag desk pipe light switch
fireplace coat rack candle bathroom stall ledge alarm clock computer tower bathroom stall cd case
kitchen counter storage organizer plunger mirror furniture music stand soap dispenser blinds backpack
toilet paper dispenser stuffed animal blackboard cart laundry detergent container toilet paper dispenser windowsill
mini fridge headphones trash can decoration dumbbell bicycle coffee table box
tray broom stair rail closet door tube light dish rack book
toaster oven guitar case box vacuum cleaner cd case clothes guitar mailbox
toilet seat cover dispenser hair dryer towel dish rack closet rod machine seat sofa chair
furniture water bottle door range hood coffee kettle furniture clock shower curtain
cart purse clothes projector screen shower head stair rail alarm clock bulletin board
storage container vent whiteboard divider keyboard piano toilet paper holder board crate
tissue box water pitcher bed bathroom counter case of water bottles floor file cabinet tube
crate bowl floor laundry hamper coat rack bucket ceiling light window
dish rack paper bag bathtub bathroom stall door folded chair stool ladder power outlet
range hood alarm clock desk ceiling light fire alarm door paper towel dispenser power strip
dustpan laundry detergent wardrobe trash bin power strip sign shower floor bathtub
handicap bar object clothes dryer bathroom cabinet calendar recycling bin stuffed animal column
mailbox ceiling light radiator structure poster shower water cooler fire alarm
music stand dumbbell shelf storage organizer luggage jacket coffee kettle storage container
bathroom counter tube potted plant bottle kitchen counter
bathroom vanity cd case mattress
laundry hamper coffee kettle
trash bin shower head
keyboard piano case of water bottles
folded chair fire alarm
luggage power strip
mattress calendar
poster
potted plant
In this supplementary material, we first present the per-phase performance of the proposed CLIMB-3D model in Appendix Appendix A, followed by the analysis of gains on rare classes achieved through the incorporation of the CBR module in Appendix Appendix B. Next, we provide detailed split information for all scenarios, based on class names, in Appendix Appendix C. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix Appendix D.
Appendix Appendix A Per-Phase Performance Analysis
We extend the analysis presented in Tab. 1 (Sec. 4.2 of the main paper) to Tab. 4 to examine the impact of our proposed method across individual splits and phases in different scenarios. The results consistently show that our method outperforms the baseline by better retaining knowledge from earlier tasks.
Split_A. In the Split_Ascenario, our model shows notable improvements across all phases. Particularly in Phase 3, although both methods show a drop in performance for s2, our approach significantly reduces forgetting, maintaining a higher $\text{mAP}_{50}$  across all splits.
Split_B. The Split_Bscenario, while more challenging due to increased semantic overlap, demonstrates our model’s strength in generalisation. In Phase 2, it achieves an overall $\text{mAP}_{50}$  of 43.13% compared to the baseline’s 24.53%. This trend continues in Phase 3, with our model achieving 31.56%, more than double the baseline’s 15.07%, indicating both improved learning and retention.
Split_C. For the Split_Cscenario, where splits are randomly grouped, Phase 1 performance is initially lower. However, in subsequent phases, our method balances stability and plasticity better than the baseline. By Phase 3, it consolidates performance across all splits, improving the overall $\text{mAP}_{50}$to 26.78%, a 5.7% gain over the baseline.
These results demonstrate the robustness of our method in handling continual learning under varied and challenging split strategies.
Appendix Appendix B Evaluation on Rare Categories
The proposed Class-Balanced Re-weighting (CBR) module, as detailed in 3.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and Pseudo-Label Generator (PLG). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
The results, shown in Tab. 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the CBR module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, show significant improvement when the CBR module is applied. Overall, the CBR module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though classes such as piano, bucket, and laundry basket are observed only once, the CBR module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+PLG module does not focus on rare classes like shower and toaster, which results in low performance, but the CBR module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed CBR module into the framework outperforms ER+PLG by 12.13%.
Appendix Appendix C Incremental Scenarios Phases
Table 7 presents the task splits for each proposed scenario introduced in Sec. 3.3, and Fig. 3 illustrates them visually for the ScanNet200 [rozenberszki2022language] dataset. The three scenarios – Split_A, Split_B, and Split_C– are each divided into three tasks: Task1, Task2, and Task3. Notably, the order of classes within these tasks is random.
Appendix Appendix D Qualitative Results
In this section, we present a qualitative comparison of the proposed framework with the baseline method. Figure 4 illustrates the results on the Split_A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach. As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline. For example, in row 1, the baseline method fails to segment the sink, while in row 3, the sofa instance is missed. Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_A  evaluation after learning all tasks.
In Figure 5, we present the results on Split_B, highlighting instances where the baseline method underperforms, marked with red circles. For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances. Similarly, in row 5, the washing machine is segmented into two instances by the baseline. In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_B  evaluation after learning all tasks.
Similarly, Figure 6 highlights the results on Split_C, where classes are encountered in random order. The comparison emphasizes the advantages of our method, as highlighted by red circles. The baseline method often misses instances or splits a single instance into multiple parts. In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_C  evaluation after learning all tasks.